# Don’t Overthink it. Preferring Shorter Thinking Chains for Improved LLM Reasoning

### v1

@@ Abstract
Reasoning large language models (LLMs) heavily rely on scaling test-time compute to perform complex reasoning tasks by generating extensive “thinking” chains. While demonstrating impressive results, this approach incurs significant computational costs and inference time. In this work, we challenge the assumption that long thinking chains results in better reasoning capabilities. We first demonstrate that shorter reasoning chains within individual questions are significantly more likely to yield correct answers—up to $34.5\%$ more accurate than the longest chain sampled for the same question. Based on these results, we suggest short-m@k, a novel reasoning LLM inference method. Our method executes $k$ independent generations in parallel and halts computation once the first $m$ thinking processes are done. The final answer is chosen using majority voting among these $m$ chains. Basic short-1@k demonstrates similar or even superior performance over standard majority voting in low-compute settings—using up to $40\%$ fewer thinking tokens. short-3@k, while slightly less efficient than short-1@k, consistently surpasses majority voting across all compute budgets, while still being substantially faster (up to $33\%$ wall time reduction). Inspired by our results, we finetune an LLM using short, long, and randomly selected reasoning chains. We then observe that training on the shorter ones leads to better performance. Our findings suggest rethinking current methods of test-time compute in reasoning LLMs, emphasizing that longer “thinking” does not necessarily translate to improved performance and can, counter-intuitively, lead to degraded results.
@@ 1 Introduction
Scaling test-time compute has been shown to be an effective strategy for improving the performance of reasoning LLMs on complex reasoning tasks (OpenAI, 2024, 2025; Team, 2025a). This method involves generating extensive thinking—very long sequences of tokens that contain enhanced reasoning trajectories, ultimately yielding more accurate solutions. Prior work has argued that longer model responses result in enhanced reasoning capabilities (Guo et al., 2025; Muennighoff et al., 2025; Anthropic, 2025). However, generating such long-sequences also leads to high computational cost and slow decoding time due to the autoregressive nature of LLMs.
@@
In this work, we demonstrate that scaling test-time compute does not necessarily improve model performance in the way previously thought. We start with a somewhat surprising observation. We take three leading reasoning LLMs, and for each generate multiple answers to each question in three complex math benchmarks. We then observe that taking the shortest answer for each question strongly and consistently outperforms both selecting a random answer (up to $18.8\%$ gap) and taking the longest one (up to $34.5\%$ gap). These performance gaps are on top of the natural reduction in sequence length—the shortest chains are $50\%$ and $67\%$ shorter than the random and longest chains respectively.
@@
Figure 1: Visual comparison between majority voting and our proposed method short-m@k with $m=1$ (“…” represent thinking time). Given $k$ parallel attempts for the same question, majority@$k$ waits until all attempts are done, and perform majority voting among them. On the other hand, our short-m@k method halts computation for all attempts as soon as the first $m$ attempts finish “thinking”, which saves much compute and time, and surprisingly also boost performance in most cases.
@@
Building on these findings, we propose short-m@k—a novel inference method for reasoning LLMs. short-m@k executes $k$ generations in parallel and terminates computation for all generations as soon as the first $m$ thinking processes are completed. The final answer is then selected via majority voting among those shortest chains, where ties are broken by taking the shortest answer among the tied candidates. A high-level, visual description of the proposed method, can be seen on Figure 1.
@@
We evaluate short-m@k using three leading reasoning LLMs, and compare it to majority voting (Wang et al., 2022)—the most common aggregation method for evaluating reasoning LLMs on complex benchmarks (Abdin et al., 2025). We show that in low-compute regimes, short-1@k, i.e., taking the single shortest chain, outperforms majority voting, while significantly reducing the time and compute needed to generate the final answer. For example, using LN-Super-$49$(Bercovich et al., 2025), short-1@k can reduce up to $40\%$ of the compute while giving the same performance as majority voting. Moreover, for high-compute regimes, short-3@k, which halts generation after three thinking chains are completed, consistently outperforms majority voting across all compute budgets, while running up to 33% faster.
@@
Based on our findings, we study whether training on short reasoning chains can lead to more accurate models. To do so, we finetune the Qwen-$2.5$-$32$B model (Yang et al., 2024) on three variants of the S1111 dataset (Muennighoff et al., 2025): S1111-short, S1111-long, and S1111-random, consisting of examples with the shortest, longest, and randomly sampled reasoning trajectories among several generations, respectively. Our experiments demonstrate that finetuning Qwen-$2.5$-$32$B on S1111-short not only yields shorter thinking lengths, but also improves model performance. Conversely, finetuning on S1111-long increases reasoning time with no significant performance gains.
@@ 2 Related work
Reasoning LLMs tackle complex tasks by employing extensive reasoning processes, often involving detailed, step-by-step trajectories (OpenAI, 2024, 2025; Team, 2025a; Abdin et al., 2025; Anthropic, 2025; Bercovich et al., 2025; Guo et al., 2025; sky, 2025; DeepMind, 2025; Team, 2025b). This capability is fundamentally based on techniques like chain-of-thought [CoT; Wei et al., 2022], which encourages models to generate intermediate reasoning steps before arriving at a final answer. By using a greater number of tokens to explore multiple problem-solving approaches, employ self-reflection, and perform verification, reasoning LLMs have demonstrated superior performance on challenging tasks such as mathematical problem-solving and code generation (Ke et al., 2025).
The ability of extensive chain-of-thought capabilities in LLMs is typically achieved through post-training methods applied to a strong base model. The two primary approaches for instilling or improving this reasoning ability are the application of reinforcement learning (RL) (Guo et al., 2025; Team, 2025a) and supervised fine-tuning (Muennighoff et al., 2025; Ye et al., 2025).
Guo et al. (2025) have demonstrated that as training progresses the model tends to generate longer thinking trajectories, which result in improved performance on complex reasoning tasks. Similarly, Anthropic (2025) and Muennighoff et al. (2025) have shown a correlation between increased average thinking length during inference and improved model performance. In this study, we challenge this assumption, demonstrating that shorter sequences are more likely to provide the correct answer.
@@
While shortening the length of CoT is beneficial for non-reasoning models (Nayab et al., 2024; Kang et al., 2025), it is higly important for reasoning LLMs as they require a very large amount of tokens to perform the thinking process. As a result, very recent studies tried to make the process more efficient, e.g., by using early exit techniques for reasoning trajectories (Pu et al., 2025; Yang et al., 2025), or by training reasoning models which enable control over the thinking length (Yu et al., 2025).
Several concurrent works studied the relationship between reasoning trajectory length and correctness. Lu et al. (2025) proposed a method for reducing the length of thinking trajectories in reasoning training datasets. Their method employs a reasoning LLM several times over an existing trajectory in order to make it shorter. As this approach eventually trains a model over shorter trajectories it is similar to the method we employ in Section 5. However, our method is simpler as it does not require an LLM to explicitly shorten the sequence. Fatemi et al. (2025) and Arora and Zanette (2025) proposed an RL methods to shorten reasoning in language models. Fatemi et al. (2025) also observed that correct answers typically require shorter thinking trajectories by averaging lengths across examples, suggesting that lengthy responses might inherently stem from RL-based optimization during training. In Section 3 we show that indeed correct answers usually use longer thinking trajectories, but also highlight that averaging across all examples might hinder this effect as easier questions require sustainably lower amount of reasoning tokens compared to harder ones. Wang et al. (2025) conducted a comprehensive study on reasoning trajectories. They found that for a specific question, correct responses from reasoning models are usually shorter than incorrect ones. We provide further analysis supporting this observation in Section 3. Moreover, our proposed inference method short-m@k is designed to enhance the efficiency of reasoning LLMs by leveraging this property.
@@ 3 Shorter thinking is preferable
As mentioned above, the common wisdom in reasoning LLMs suggests that increased test-time computation enhances model performance. Specifically, it is widely assumed that longer reasoning process, which entails extensive reasoning thinking chains, correlates with improved task performance (OpenAI, 2024; Anthropic, 2025; Muennighoff et al., 2025). We challenge this assumption and ask whether generating more tokens actually leads to better performance. To that end, we generate multiple answers per question and compare performance based solely on the shortest, longest and randomly sampled thinking chains among the generated samples.
@@ 3.1 Experimental details
We consider three leading, high-performing, open, reasoning LLMs. Llama-$3.3$-Nemotron-Super-$49$B-v1111 [LN-Super-$49$B; Bercovich et al., 2025]: a reasoning RL-enhanced version of Llama-$3.3$-$70$(Grattafiori et al., 2024); R1111-Distill-Qwen-$32$B [R1111-$32$B; Guo et al., 2025]: an SFT finetuned version of Qwen-$2.5$-$32$B-Instruct (Yang et al., 2024) derived from R1111 trajectories; and QwQ-$32$B a reasoning RL-enhanced version Qwen-$2.5$-$32$B-Instruct (Team, 2025a).
We evaluate all models using three competitive reasoning benchmarks: AIME $2024$ (of America, 2024), AIME $2025$ (of America, 2025) and HMMT February $2025$, from the Math Arena benchmark (Balunović et al., 2025). The three benchmarks are derived from math competitions, and involve solving problems that cover a broad range of mathematics topics. Each dataset consists of $30$ examples with varied difficulty.
For each question, we generate $20$ responses per model, yielding a total of $5$,$400$ generations. For all models we use temperature of $0.7$, top-p=$0.95$ and a maximum number of generated tokens of $32$,$768$. When measuring the thinking chain length, we measure the token count between the <think> and </think> tokens. We run inference for all models using paged attention via the vLLM framework (Kwon et al., 2023). [^1] We used $8$ H100100100100 GPUs per model. The total decoding time for all models sums up to about $110$ hours.
@@ 3.2 The shorter the better
We first note that as observed in recent studies (Anthropic, 2025; OpenAI, 2024; Muennighoff et al., 2025), thinking chains tend to be longer for harder questions. To quantify this phenomenon with our generated samples, for each model we split the questions into three equal size groups according to model’s success rate. Then, we calculate the average thinking length per easier and harder questions. [^2] In this section we exclude generations where thinking is not completed within the maximum generation length, as these often result in an infinite thinking loop. We also provide the average lengths for the correct and incorrect attempts per split.
Table 1: Average thinking tokens for correct (C), incorrect (IC) and all (A) answers, per easier and harder questions. The numbers are in thousands of tokens.
Model Easy Hard
C/IC/A C/IC/A
LN-Super-49B $\phantom{0}5.3/11.1/5.7$ $12.4/16.8/16.6$
R1-32B $\phantom{0}4.9/13.7/5.3$ $14.4/15.8/15.7$
QwQ-32B [^3] The QwQ-32B model correctly answered all of its easier questions in all attempts. 8.4/$\phantom{0.}\text{--}\phantom{0}/8.4$ – / 8.4 $19.1/22.8/22.3$
Table 1 shows that indeed models use more tokens for more challenging questions, up to a factor of $2.9$. This may lead to the assumption that longer thinking chains leads to more complex reasoning, and therefore, better performance. Nevertheless, surprisingly, we also observe that, within each question subset, correct answers are typically shorter than incorrect ones. To study the connection between performance and thinking length in a controlled manner, we turn to compare different answers to the same question.
We compare short vs. long thinking chains for the same question, along with a random chain. Results are presented in Table 2. First, as expected, the shortest answers are $25\%$$50\%$ shorter compared to randomly sampled responses. However, we also note that across almost all models and benchmarks, considering the answer with the shortest thinking chain actually boosts performance, yielding an average absolute improvement of $2.2\%$$15.7\%$ across benchmarks compared to randomly selected generations. When considering the longest thinking answers among the generations, we further observe an increase in thinking chain length, with up to $75\%$ more tokens per chain. These extended reasoning trajectories substantially degrade performance, resulting in average absolute reductions ranging between $12\%$$18.8\%$ compared to random generations. These trends are most noticeable when comparing the shortest generation with the longest ones, with an absolute performance gain of up to $34.5\%$ in average accuracy and a substantial drop in the number of thinking tokens.
The above results suggest that long generations might come with a significant price-tag, not only in running time, but also in performance. While more complex questions generally require a greater number of thinking tokens, within an individual example, shorter thinking trajectories are much more likely to be correct. This observation challenges the assumption for reasoning LLMs that longer reasoning processes typically improve performance. Next, we propose strategies to leverage these findings to improve the efficiency and effectiveness of reasoning LLMs.
@@
AIME 2024 AIME 2025 HMMT Average
Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$
@@
random 11258 58.8 12105 51.3 13445 33.0 12270 47.7
longest 18566 33.3 18937 30.0 19790 23.3 19098 $(+56\%)$ 28.9
shortest 06276 76.7 07036 66.7 07938 46.7 7083 $(-42\%)$ 63.4
@@
random 9614 71.8 11558 56.4 12482 38.3 11218 55.5
longest 17689 53.3 19883 36.7 20126 23.3 19233 $(+71\%)$ 37.8
shortest 04562 80.0 06253 63.3 06557 36.7 5791 $(-48\%)$ 60.0
@@
random 13093 82.0 14495 72.3 16466 52.5 14685 68.9
longest 20059 70.0 21278 63.3 24265 36.7 21867 $(+49\%)$ 56.7
shortest 08655 86.7 10303 66.7 11370 60.0 10109 $(-31\%)$ 71.1
@@
Based on the results presented in Section 3, we suggest a novel inference method for reasoning LLMs. Our method—short-m@k—leverages batch inference of LLMs per question, using multiple parallel decoding runs for the same query. We begin by introducing our method in Section 4.1. We then describe our evaluation methodology, which takes into account inference compute and running time (Section 4.2). Finally, we present our results (Section 4.3) and ablation studies (Section 4.4).

### v2

@@ Abstract
Reasoning large language models (LLMs) heavily rely on scaling test-time compute to perform complex reasoning tasks by generating extensive “thinking” chains. While demonstrating impressive results, this approach incurs significant computational costs and inference time. In this work, we challenge the assumption that long thinking chains results in better reasoning capabilities. We first demonstrate that shorter reasoning chains within individual questions are significantly more likely to yield correct answers—up to $34.5\%$ more accurate than the longest chain sampled for the same question. Based on these results, we suggest short-m@k, a novel reasoning LLM inference method. Our method executes $k$ independent generations in parallel and halts computation once the first $m$ thinking processes are done. The final answer is chosen using majority voting among these $m$ chains. Basic short-1@k demonstrates similar or even superior performance over standard majority voting in low-compute settings—using up to $40\%$ fewer thinking tokens. short-3@k, while slightly less efficient than short-1@k, consistently surpasses majority voting across all compute budgets, while still being substantially faster (up to $33\%$ wall time reduction). To further validate our findings, we finetune LLMs using short, long, and randomly selected reasoning chains. We then observe that training on the shorter ones leads to better performance. Our findings suggest rethinking current methods of test-time compute in reasoning LLMs, emphasizing that longer “thinking” does not necessarily translate to improved performance and can, counter-intuitively, lead to degraded results.
@@ 1 Introduction
Scaling test-time compute has been shown to be an effective strategy for improving the performance of reasoning LLMs on complex reasoning tasks (OpenAI, 2024; 2025; Team, 2025b). This method involves generating extensive thinking—very long sequences of tokens that contain enhanced reasoning trajectories, ultimately yielding more accurate solutions. Prior work has argued that longer model responses result in enhanced reasoning capabilities (Guo et al., 2025; Muennighoff et al., 2025; Anthropic, 2025). However, generating such long-sequences also leads to high computational cost and slow decoding time due to the autoregressive nature of LLMs.
@@
In this work, we demonstrate that scaling test-time compute does not necessarily improve model performance as previously thought. We start with a somewhat surprising observation. We take four leading reasoning LLMs, and for each generate multiple answers to each question in four complex reasoning benchmarks. We then observe that taking the shortest answer for each question strongly and consistently outperforms both a strategy that selects a random answer (up to $18.8\%$ gap) and one that selects the longest answer (up to $34.5\%$ gap). These performance gaps are on top of the natural reduction in sequence length—the shortest chains are $50\%$ and $67\%$ shorter than the random and longest chains, respectively.
@@
Figure 1: Visual comparison between majority voting and our proposed method short-m@k with $m=1$ (“…” represent thinking time). Given $k$ parallel attempts for the same question, majority@$k$ waits until all attempts are done, and perform majority voting among them. On the other hand, our short-m@k method halts computation for all attempts as soon as the first $m$ attempts finish “thinking”, which saves compute and time, and surprisingly also boost performance in most cases.
@@
Building on these findings, we propose short-m@k—a novel inference method for reasoning LLMs. short-m@k executes $k$ generations in parallel and terminates computation for all generations as soon as the first $m$ thinking processes are completed. The final answer is then selected via majority voting among those shortest chains, where ties are broken by taking the shortest answer among the tied candidates. See Figure˜1 for visualization.
@@
We evaluate short-m@k using six reasoning LLMs, and compare it to majority voting—the most common aggregation method for evaluating reasoning LLMs on complex benchmarks (Wang et al., 2022; Abdin et al., 2025). We show that in low-compute regimes, short-1@k, i.e., taking the single shortest chain, outperforms majority voting, while significantly reducing the time and compute needed to generate the final answer. For example, using LN-Super-4949(Bercovich and others, 2025), short-1@k can reduce up to $40\%$ of the compute while giving the same performance as majority voting. Moreover, for high-compute regimes, short-3@k, which halts generation after three thinking chains are completed, consistently outperforms majority voting across all compute budgets, while running up to $33\%$ faster.
To gain further insights into the underlining mechanism of why shorter thinking is preferable, we analyze the generated reasoning chains. We first show that while taking the shorter reasoning is beneficial per individual question, longer reasoning is still needed to solve harder questions, as claimed in recent studies (Anthropic, 2025; OpenAI, 2024; Muennighoff et al., 2025). Next, we analyze the backtracking and re-thinking behaviors of reasoning chains. We find that shorter reasoning paths are more effective, as they involve fewer backtracks, with a longer average backtrack length. This finding holds both generally and when controlling for overall trajectory length.
@@
To further strengthening our findings, we study whether training on short reasoning chains can lead to more accurate models. To do so, we finetune two Qwen-$2.5$ (Yang and others, 2024) models (77B and 3232B) on three variants of the S11 dataset (Muennighoff et al., 2025): S11-short, S11-long, and S11-random, consisting of examples with the shortest, longest, and randomly sampled reasoning trajectories among several generations, respectively. Our experiments demonstrate that finetuning using S11-short not only yields shorter thinking lengths, but also improves model performance. Conversely, finetuning on S11-long increases reasoning time with no significant performance gains.
@@ 2 Related work
Reasoning LLMs tackle complex tasks by employing extensive reasoning processes, often involving detailed, step-by-step trajectories (OpenAI (2024); OpenAI (2025); Q. Team (2025b); M. Abdin, S. Agarwal, A. Awadallah, V. Balachandran, H. Behl, L. Chen, G. de Rosa, S. Gunasekar, M. Javaheripi, N. Joshi, et al. (2025); Anthropic (2025); A. Bercovich et al. (2025); D. Guo, D. Yang, H. Zhang, J. Song, R. Zhang, R. Xu, Q. Zhu, S. Ma, P. Wang, X. Bi, et al. (2025); 27; G. DeepMind (2025); Q. Team (2025a)). This capability is fundamentally based on techniques like chain-of-thought (CoT; Wei et al., 2022), which encourage models to generate intermediate reasoning steps before arriving at a final answer. Modern LLMs use a large number of tokens, often referred to as “thinking tokens”, to explore multiple problem-solving approaches, to employ self-reflection, and to perform verification. This thinking capability has allowed them to achieve superior performance on challenging tasks such as mathematical problem-solving and code generation (Ke et al., 2025).
The LLM thinking capability is typically achieved through post-training methods applied to a strong base model. The two primary approaches to instilling or improving this reasoning ability are using reinforcement learning (RL) (Guo et al., 2025; Team, 2025b) and supervised fine-tuning (Muennighoff et al., 2025; Ye et al., 2025). Guo et al. (2025) have demonstrated that as training progresses the model tends to generate longer thinking trajectories, which results in improved performance on complex tasks. Similarly, Anthropic (2025) and Muennighoff et al. (2025) have shown a correlation between increased average thinking length during inference and improved performance. We challenge this assumption, demonstrating that shorter sequences are more likely to yield an accurate answer.
@@
While shortening the length of CoT is beneficial for non-reasoning models (Nayab et al., 2024; Kang et al., 2025), it is higly important for reasoning LLMs as they require a very large amount of tokens to perform the thinking process. As a result, recent studies tried to make the process more efficient, e.g., by using early exit techniques for reasoning trajectories (Pu et al., 2025; Yang et al., 2025), by suppressing backtracks (Wang et al., 2025a) or by training reasoning models which enable control over the thinking length (Yu et al., 2025).
Several recent works studied the relationship between reasoning trajectory length and correctness. Lu et al. (2025) proposed a method for reducing the length of thinking trajectories in reasoning training datasets. Their method employs a reasoning LLM several times over an existing trajectory in order to make it shorter. As this approach eventually trains a model over shorter trajectories it is similar to the method we employ in Section˜6. However, our method is simpler as it does not require an LLM to explicitly shorten the sequence. Fatemi et al. (2025); Qi et al. (2025) and Arora and Zanette (2025) proposed RL methods to shorten reasoning in language models. Fatemi et al. (2025) also observed that correct answers typically require shorter thinking trajectories by averaging lengths across examples, suggesting that lengthy responses might inherently stem from RL-based optimization during training. In Section˜5.1 we show that indeed correct answers usually use shorter thinking trajectories, but also highlight that averaging across all examples might hinder this effect as easier questions require sustainably lower amount of reasoning tokens compared to harder ones.
More relevant to our work, Wu et al. (2025) showed that there is an optimal thinking length range for correct answers according to the difficulty of the question, while Wang et al. (2025b) found that for a specific question, correct responses from reasoning models are usually shorter than incorrect ones. We provide further analysis supporting these observations in Sections˜3 and 5. Finally, our proposed inference method short-m@k is designed to enhance the efficiency of reasoning LLMs by leveraging this property, which can be seen as a generalization of the FFS method (Agarwal et al., 2025), which selects the shortest answer among several candidates as in our short-1@k.
@@ 3 Shorter thinking is preferable
As mentioned above, the common wisdom in reasoning LLMs suggests that increased test-time computation enhances model performance. Specifically, it is widely assumed that longer reasoning process, which entails extensive reasoning thinking chains, correlates with improved task performance (OpenAI, 2024; Anthropic, 2025; Muennighoff et al., 2025). We challenge this assumption and ask whether generating more tokens per trajectory actually leads to better performance. To that end, we generate multiple answers per question and compare performance based solely on the shortest, longest and randomly sampled thinking chains among the generated samples.
@@ 3.1 Experimental details
We consider four leading, high-performing, open, reasoning LLMs. Llama-$3.3$-Nemotron-Super-4949B-v11 [LN-Super-4949B; Bercovich and others, 2025]: a reasoning RL-enhanced version of Llama-3.33.3-7070(Grattafiori et al., 2024); R11-Distill-Qwen-3232B [R11-3232B; Guo et al., 2025]: an SFT finetuned version of Qwen-2.52.5-3232B-Instruct (Yang and others, 2024) derived from R11 trajectories; QwQ-3232B a reasoning RL-enhanced version Qwen-2.52.5-3232B-Instruct (Team, 2025b); and R1-$0528$ a 670670B RL-trained flagship reasoning model (R11-670670B; Guo et al., 2025). We also include results for smaller models in Appendix˜D.
We evaluate all models using four competitive reasoning benchmarks. We use AIME $2024$ (of America, 2024), AIME $2025$ (of America, 2025) and HMMT February $2025$, from the Math Arena benchmark (Balunović et al., 2025). This three benchmarks are derived from math competitions, and involve solving problems that cover a broad range of mathematics topics. Each dataset consists of $30$ examples with varied difficulty. We also evaluate the models using the GPQA-diamond benchmark [GPQA-D; Rein et al., 2024], which consists of $198$ multiple-choice scientific questions, and is considered to be challenging for reasoning LLMs (DeepMind, 2025).
For each question, we generate $20$ responses per model, yielding a total of about 3636k generations. For all models we use temperature of $0.7$, top-p=$0.95$ and a maximum number of generated tokens of 3232,$768$. When measuring the thinking chain length, we measure the token count between the <think> and </think> tokens. We run inference for all models using paged attention via the vLLM framework (Kwon et al., 2023).
@@ 3.2 The shorter the better
GPQA-D AIME 2024 AIME 2025 HMMT Math Average
Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$
@@
random 5357 65.1 11258 58.8 12105 51.3 13445 33.0 12270 47.7
longest 8763 $(+64\%)$ 57.6 18566 33.3 18937 30.0 19790 23.3 19098 $(+56\%)$ 28.9
shortest 2790 $(-48\%)$ 69.1 06276 76.7 07036 66.7 07938 46.7 7083 $(-42\%)$ 63.4
@@
random 5851 62.5 9614 71.8 11558 56.4 12482 38.3 11218 55.5
longest 9601 $(+64\%)$ 57.1 17689 53.3 19883 36.7 20126 23.3 19233 $(+71\%)$ 37.8
shortest 3245 $(-45\%)$ 64.7 04562 80.0 06253 63.3 06557 36.7 5791 $(-48\%)$ 60.0
@@
random 8532 63.7 13093 82.0 14495 72.3 16466 52.5 14685 68.9
longest 12881 $(+51\%)$ 54.5 20059 70.0 21278 63.3 24265 36.7 21867 $(+49\%)$ 56.7
shortest 5173 $(-39\%)$ 64.7 08655 86.7 10303 66.7 11370 60.0 10109 $(-31\%)$ 71.1
R1-670B
random 11843 76.2 16862 83.8 18557 82.5 21444 68.2 18954 78.2
longest 17963 $(+52\%)$ 63.1 22603 70.0 23570 66.7 27670 40.0 24615 $(+30\%)$ 58.9
shortest 8116 $(-31\%)$ 75.8 11229 83.3 13244 83.3 13777 83.3 12750 $(-33\%)$ 83.3
For all generated answers, we compare short vs. long thinking chains for the same question, along with a random chain. Results are presented in Table˜1.111In this section we exclude generations where thinking is not completed within the maximum generation length, as these often result in an infinite thinking loop. First, as expected, the shortest answers are $25\%$$50\%$ shorter compared to randomly sampled responses. However, we also note that across almost all models and benchmarks, considering the answer with the shortest thinking chain actually boosts performance, yielding an average absolute improvement of $2.2\%$$15.7\%$ on the math benchmarks compared to randomly selected generations. When considering the longest thinking answers among the generations, we further observe an increase in thinking chain length, with up to $75\%$ more tokens per chain. These extended reasoning trajectories substantially degrade performance, resulting in average absolute reductions ranging between $5.4\%$$18.8\%$ compared to random generations over all benchmarks. These trends are most noticeable when comparing the shortest generation with the longest ones, with an absolute performance gain of up to $34.5\%$ in average accuracy and a substantial drop in the number of thinking tokens.
The above results suggest that long generations might come with a significant price-tag, not only in running time, but also in performance. That is, within an individual example, shorter thinking trajectories are much more likely to be correct. In Section˜5.1 we examine how these results relate to the common assumption that longer trajectories leads to better LLM performance. Next, we propose strategies to leverage these findings to improve the efficiency and effectiveness of reasoning LLMs.
@@
Based on the results presented in Section˜3, we suggest a novel inference method for reasoning LLMs. Our method—short-m@k—leverages batch inference of LLMs per question, using multiple parallel decoding runs for the same query. We begin by introducing our method in Section˜4.1. We then describe our evaluation methodology, which takes into account inference compute and running time (Section˜4.2). Finally, we present our results (Section˜4.3).